### v2

# Entropy production of the contact model

Abstract
We propose an expression for the production of entropy for system described by a stochastic dynamics which is appropriate for the case where the reverse transition rate vanishes but the forward transition is nonzero. The expression is positive definite and based on the inequality $x\ln x-(x-1)\geq 0$. The corresponding entropy flux is linear in the probability distribution allowing its calculation as an average. The expression is applied to the one-dimensional contact process at the stationary state. We found that the rate of entropy production per site is finite with a singularity at the critical point with diverging slope.
I Introduction
The entropy production is the central concept of the theory of systems out of thermodynamic equilibirum. This quantity characterizes the state of nonequilibrium and is nonzero not only in systems that have not yet reached thermodynamic equilibrium but also in systems that are in a permanent nonequilibrium stationary state. Within the stochastic approach to thermodynamics the production of entropy is related to the ratio between the probability of a trajectory and its reverse.
Let us consider a stochastic dynamics in continuous time and discrete space of states and let $P_{i}$ be the probability distribution of states $i$, and $W_{ij}$ be the transition probability rate from state $j$ to state $i$. The equation governing the time evolution of $P_{i}$ is assumed to be given by a master equation kampen1981 ; tome2015L
$\frac{dP_{i}}{dt}=\sum_{j}(W_{ij}P_{j}-W_{ji}P_{i}).$ (1)
The total production of entropy is a sum of terms each one corresponding to a transition $j\to i$, and given by the Schnakenberg expression schnakenberg1976
$\Pi_{ij}=\frac{1}{2}(W_{ij}P_{j}-W_{ji}P_{i})\ln\frac{W_{ij}P_{j}}{W_{ji}P_{i}}.$ (2)
An identical expression is valid for the reverse transition $i\to j$. The expression (2) is clearly nonnegative because if we let $W_{ij}P_{j}/W_{ji}P_{i}=x$ then this expression is proportional to
$(x-1)\ln x\geq 0,$ (3)
which is always nonnegative, reflecting the main feature of the production of entropy. This is the usual expression for the rate of entropy production and has been extensively studied and applied luo1984 ; mou1986 ; lebowitz1999 ; maes2003 ; seifert2005 ; tome2006 ; zia2007 ; andrieux2007 ; harris2007 ; gaveau2009 ; esposito2009 ; szabo2010 ; tome2010 ; hinrichsen2011 ; esposito2012 ; tome2015 ; tome2018 . In particular it has been used to determined the rate of entropy production of irreversible stochastic lattice models crochik2005 ; tome2012 ; noa2019 ; tome2023 ; hawthorne2023 at the nonequilibrium stationary state.
The expression (2) however breaks down when one transition rate vanishes while the other remains nonzero, that is, when a transition has no reverse. There are many examples of stochastic systems where this happens, the most striking examples being the systems with absorbing states such as the contact process marro1999 ; hinrichsen2000 ; henkel2008 . In this model, particles residing on a lattice are annihilated and created according to the following rules. A particle is annihilated spontaneously, and is created in a empty site if this site has at least one neighboring site occupied. Thus the transition corresponding to the annihilation of an isolated particle has no reverse, and the application of the formula (2) would result in a divergent rate of entropy production.
It should be pointed out that the Schnakenberg formula has a straight relationship with the Jarzynski equality. Thus the breakdown of this formula results in the inapplicability of the Jarzynski inequality when the probability of the backward probability vanishes wereas the forward probability remains finite murashita2014 .
This problem has been addressed by several authors and there have been some attempts to overcome it murashita2014 ; ohkubo2009 ; benavraham2011 ; zeraati2012 ; rahav2014 ; saha2016 ; pal2017 ; busiello2020 ; pal2021 ; pal2021a . Some of these attempts focused mainly on the introduction of an effective backward transitions while keeping formula (2) unchanged. A distinguishing formula was proposed by Busiello, Gupta, and Maritan busiello2020 and will be commented further on. Here we take a different route. To determine the rate of entropy production when the reverse transition is absent, that is, when $W_{ij}>0$ but $W_{ji}=0$, we propose the following expression
$\Pi_{ij}=W_{ij}P_{j}\ln\frac{P_{j}}{P_{i}}-W_{ij}(P_{j}-P_{i}),$ (4)
which is associated to the transition $j\to i$. The production associated to the reverse transition $i\to j$ is defined as being zero. We see that the expression (4) is clearly nonnegative because if we let $P_{j}/P_{i}=x$ then this expression is proportional to
$x\ln x-(x-1)\geq 0.$ (5)
The present approach is applied to the calculation of the rate of entropy production of the one-dimensional contact process which is known to undergo a phase transition from an active to an absorbing state as one increases the annihilation rate constant. The rate of entropy production is found to be finite for any value of the annihilation parameter displaying a singularity at the critical point.
II Entropy production
The entropy $S$ of a system described by the master equation (1) is given by the Gibbs expression
$S=-\sum_{i}P_{i}\ln P_{i}.$ (6)
Deriving $S$ with respect to time we get
$\frac{dS}{dt}=-\sum_{i}\frac{dP_{i}}{dt}\ln P_{i}.$ (7)
Using the master equation (1) we reach the expression
$\frac{dS}{dt}=\sum_{ij}W_{ij}P_{j}\ln\frac{P_{j}}{P_{i}},$ (8)
which is well defined no matter if $W_{ij}$ vanishes or not.
The total rate of entropy production $\Pi$ is understood as a sum of terms $\Pi_{ij}$,
$\Pi=\sum_{ij}\Pi_{ij},$ (9)
each one associated to a transition $j\to i$ and defined as follows. If $W_{ij}$ and its reverse $W_{ji}$ are nonzero, $\Pi_{ij}$ is given by the expression (2). If $W_{ij}$ is nonzero but the reverse transition rate $W_{ij}$ vanishes, it is given by the expression (4).
If we add $dS/dt$ and $\Pi$, we find the total entropy flux $\Psi$, that is, $\Psi=dS/dt+\Pi$, or
$\frac{dS}{dt}=\Pi-\Psi,$ (10)
that is, the time variation of the entropy of a system is equal to the rate of entropy production minus the flux of entropy from the system to the outside.
Writing $\Psi$ as a sum of terms
$\Psi=\sum_{ij}\Psi_{ij},$ (11)
we get the expressions for $\Psi_{ij}$, understood as the flux of entropy associated to a transition $j\to i$. It is given by
$\Psi_{ij}=\frac{1}{2}(W_{ij}P_{j}-W_{ji}P_{i})\ln\frac{W_{ij}}{W_{ji}},$ (12)
if $W_{ij}$ and its reverse $W_{ji}$ are nonzero, and by
$\Psi_{ij}=-W_{ij}(P_{j}-P_{i}),$ (13)
if $W_{ij}$ is nonzero but the reverse transition $W_{ij}$ vanishes. In this case we set $\Psi_{ji}=0$. Both expressions (12) and (13) are linear in $P_{i}$ so that the entropy flux can be understood as an average over the distribution of probability $P_{i}$.
In the stationary state $dS/dt$ vanishes and $\Pi=\Psi$. If the stationary state is a nonequilibrium state then $\Pi=\Psi>0$ and entropy are constantly being produced and thrown away into the environment as $\Psi>0$.
III Stochastic lattice models
Stochastic lattice models are defined on a lattice where each site $i$ holds a stochastic variable $\sigma_{i}$ that takes discrete values. We consider the simplest case where it takes only two values, 0 and 1, interpreted as the site being empty or occupied by a particle. We denote by $\overline{\sigma}_{i}$ the variable related to $\sigma_{i}$ in such a way that when the latter takes one of the two values, the former takes the other. The whole state of the system is given by all $\sigma_{i}$ and is denoted by $\sigma$, that is,
$\sigma=(\sigma_{1},\sigma_{2},\ldots,\sigma_{i},\ldots).$ (14)
It is convenient to define the state $\sigma^{i}$, by
$\sigma^{i}=(\sigma_{1},\sigma_{2},\ldots,\overline{\sigma}_{i},\ldots).$ (15)
We restrict ourselves to models that changes only the state of one site. The transition probability of changing from $\sigma$ to $\sigma^{i}$ is denoted by $w_{i}(\sigma)$ and the master equation that governs the time evolution of the probability distribution $P(\sigma)$ is given by
$\frac{d}{dt}P(\sigma)=\sum_{i}\{w_{i}(\sigma^{i})P(\sigma^{i})-w_{i}(\sigma)P(\sigma)\}.$ (16)
The entropy is given by
$S=-\sum_{\sigma}P(\sigma)\ln P(\sigma),$ (17)
and its time variation is
$\frac{dS}{dt}=\sum_{i}\sum_{\sigma}w_{i}(\sigma)P(\sigma)\ln\frac{P(\sigma)}{P(\sigma^{i})},$ (18)
The rate of entropy production $\Pi$ is a sum of terms,
$\Pi=\sum_{i}\Pi_{i},$ (19)
each one given by
$\Pi_{i}=\frac{1}{2}\sum_{\sigma}\{w_{i}(\sigma)P(\sigma)-w_{i}(\sigma^{i})P(\sigma^{i})\}\ln\frac{w_{i}(\sigma)P(\sigma)}{w_{i}(\sigma^{i})P(\sigma^{i})},$ (20)
if $w_{i}(\sigma)$ and $w_{i}(\sigma^{i})$ are both nonzero, and
$\Pi_{i}=\sum_{\sigma}w_{i}(\sigma)\{P(\sigma)\ln\frac{P(\sigma)}{P(\sigma^{i})}-P(\sigma)+P(\sigma^{i})\},$ (21)
if $w_{i}(\sigma)$ is nonzero but $w_{i}(\sigma^{i})$ vanishes.
In an analogous manner, the entropy flux $\Psi$ is a sum of terms,
$\Psi=\sum_{i}\Psi_{i},$ (22)
each one given by
$\Psi_{i}=\frac{1}{2}\sum_{\sigma}\{w_{i}(\sigma)P(\sigma)-w_{i}(\sigma^{i})P(\sigma^{i})\}\ln\frac{w_{i}(\sigma)}{w_{i}(\sigma^{i})},$ (23)
if $w_{i}(\sigma)$ and $w_{i}(\sigma^{i})$ are both nonzero, and
$\Psi_{i}=-\sum_{\sigma}w_{i}(\sigma)\{P(\sigma)-P(\sigma^{i})\},$ (24)
if $w_{i}(\sigma)$ is nonzero but $w_{i}(\sigma^{i})$ vanishes.
IV Contact model
We apply the results above to a specific stochastic lattice model, namely, the one-dimensional contact process. The model is defined on a chain where each site can be occupied by a particle or empty. The probability of creating a particle in an empty site is equal to the fraction of occupied neighboring sites, and the probability of annihilation of a particle is $p$. Below we write down the possible transitions rates related to a site and its two nearest neighbors
$\begin{array}[]{lll}101\to 111&1&\\ 100\to 110&1/2&\\ 001\to 011&1/2&\\ 000\to 010&0&\end{array}\hskip 28.45274pt\begin{array}[]{lll}111\to 101&p&\\ 110\to 100&p&\\ 011\to 001&p&\\ 010\to 000&p&\end{array}$
Using the equations (23) and (24), we write down the entropy flux associated to the above transitions, that are related to three given consecutive sites of the lattice. For the transitions $101\to 111$ and $111\to 101$, the fluxes are the same and given by
$\psi_{1}=\frac{1}{2}(P_{101}-pP_{111})\ln\frac{1}{p},$ (25)
where $P_{101}$ and $P_{111}$ are the probabilities of three consecutive sites being in states 101 and 111, respectively. The contribution coming from the transitions $100\to 110$ and $110\to 100$ are equal and given by
$\psi_{2}=\frac{1}{2}(\frac{1}{2}P_{100}-pP_{110})\ln\frac{1/2}{p}$ (26)
whereas that associated to the transitions $001\to 011$ and $011\to 001$ are equal and given by
$\psi_{3}=\frac{1}{2}(\frac{1}{2}P_{001}-pP_{011})\ln\frac{1/2}{p}.$ (27)
The contribution coming from $010\to 000$ is
$\psi_{4}=-p(P_{010}-P_{000}).$ (28)
We point out that this last expression is valid when $P_{010}$ is nonzero. If $P_{010}$ vanishes then $\psi_{4}$ should be set equal to zero even if $P_{000}$ is nonzero. The total entropy flux associated to a triple of consecutive sites is $\psi=2\psi_{1}+2\psi_{2}+2\psi_{3}+\psi_{4}$.
We have performed numerical simulations on a chain with periodic boundary conditions and a number of sites $L$ ranging from 100 to 2000. The simulation was carried out as follows tome2015L . At each time step, a site is chosen at random and it is updated according to the rules above. Our simulations is performed in such a way that we do not allow the last particle to be annihilated tome2015L . Thus strictly speaking there is no absorbing states. However, in the thermodynamic limit the system may reach a state with vanishing density, which characterizes the absorbing state. This indeed happens if the parameter $p$ is above its critical value $p_{c}$. That is, in the stationary state, and in the thermodynamic limit, the system is found in the state such that for $p\geq p_{c}$, the density of particle $\rho$ vanishes whereas for $p<p_{c}$, the system is in an active state with $\rho\neq 0$.
Figure 1: Entropy production rate $\psi$ as a function of the annihilation parameter $p$, obtained from simulations on chains of several sizes. The vertical dashed line indicates the critical value $p_{c}=0.303228$. The inset shows an enlargement around the critical point.
The eight three-site probabilities that appear in the epressions (25), (26), (27), and (28) are determined from the numerical simulations from which we get $\psi_{1}$, $\psi_{2}$, $\psi_{3}$, $\psi_{4}$, and $\psi$. We recall that $\psi$ is also the rate of entropy production, because we are considering the stationary state. In figure 1 we show $\psi$ for several values of $L$. As one increases the parameter $p$, the flux of entropy per site $\psi$ also increases, reaches an inflection point, and then decreases only after reaching a maximum. The inflection point occurs near the critical point that takes place at $p_{c}=0.303228$ marro1999 . From the inset of figure 1 we see that the slope at the inflection point increases with the size of the lattice.
These results indicates that in the thermodynamic limit, $\psi$ is finite at the critical point with a diverging slope at this point. We thus assume that the singularity is described by
$\psi_{0}-\psi\sim\varepsilon^{b},$ (29)
where $\varepsilon=p_{c}-p$ is the deviation of the parameter $p$ from its critical value, and $10<b<$ 1. We remark that for $p>p_{c}$, the entropy flux per site $\psi$ vanishes as $L\to\infty$ as can be inferred from the figure 1, and because in this limit the density vanishes. The singularity of the derivative $\Gamma=d\psi/dp$ is described by
$\Gamma\sim\varepsilon^{-a},$ (30)
where $a=1-b$, and diverges at the critical point, as can seen in figure 2.
Using a finite size scaling we write
$\Gamma_{L}=L^{a/\nu}f(L\varepsilon^{\nu}),$ (31)
where $f(x)$ is a scaling function, and $\nu$ is the critical exponent associated to the correlation length. From this scaling it follows that the slope at the critical point increases with $L$ with the following behavior
$\ln\Gamma_{L}=\frac{a}{\nu}\ln L.$ (32)
From the numerical values of the slopes of $\psi$, we obtain the following value $a/\nu=0.67\pm 0.05$. Using the value $\nu=1.097$ marro1999 for the one-dimensional contact model we find $a=0.73\pm 0.05$, and $b=0.27\pm 0.02$.
The results above indicates that both the production of entropy $\psi$ and its derivative $\Gamma=d\psi/dp$ have a singular behavior at the critical point that is described by a power law with exponents $b$ and $a$. We point out that the numerical value of the exponent $b$ for the production of entropy is numerically equal to the exponent $\beta$ associated to the order parameter $\rho$ of the one-dimensional contact model which is $\beta=0.2765$ marro1999 . This result comes from the fact that the entropy production is dominated by $P(010)$ which has a critical behavior similar to the density of particles which is the order parameter.
Figure 2: Derivative of the entropy production rate $\Gamma=d\psi/dp$ as a function of the annihilation parameter $p$, obtained from the data of figure 1. The vertical dashed line indicates the critical value $p_{c}=0.303228$. The inset shows a logarithm plot of $\Gamma$ around the critical point.
V Unidirectional transition rates
Let us consider that the states 1 and 2 are connected by a unidirectional transition rate $W_{21}>0$ from 1 to 2. The contribution of this transition rate to $dP_{1}/dt$ is $-W_{21}P_{1}$ and to $dP_{2}/dt$ is $W_{21}P_{1}$, and the contribubtion to $dS/dt$ is
$W_{21}P_{1}\ln\frac{P_{1}}{P_{2}}.$ (33)
Let us consider the case of a system such that at the stationary state the difference $P_{1}-P_{2}$ is small. In this case the expression above becomes
$-W_{21}(P_{2}-P_{1}).$
At the stationary state $P_{1}$ and $P_{2}$ does not depend on time and $P_{2}\geq P_{1}$, and this contribution to $dS/dt$ is negative. Since at the stationary state $dS/dt$ vanishes then there should be positive contributions coming from the bidirectional transitions. If the transitions are all unidirectional then at the stationary state $P_{1}=P_{2}$ and all probabilities are equal. Taking into account that $dS/dt$ is equal to the entropy production $\Pi$ minus the flux of entropy $\Psi$ then the term above should be a part of the flux of entropy, that is, the contribution to $\Psi$ coming from the unidirectional transition is
$\Psi_{21}=W_{21}(P_{2}-P_{1}).$ (34)
Summing this term with (33), we find the contribution of the unidirectional transition to the rate of entropy production
$\Pi_{21}=W_{21}P_{1}\ln\frac{P_{1}}{P_{2}}+W_{21}(P_{2}-P_{1}).$ (35)
which turns out to be positive definite, and is our proposed expression (4).
The above reasoning does not show that the form (34) is necessary but shows that it is sufficient. In fact any other form $F(P_{2},P_{1})$ such that $F\geq 0$ for $P_{2}\geq P_{1}$ will serve as long as
$P_{1}\ln\frac{P_{1}}{P_{2}}+F(P_{2},P_{1})\geq 0,$ (36)
for any $P_{1}$ and $P_{2}$. Writing $F(P_{2},P_{1})=P_{1}f(P_{2}/P_{1})$ this condition becomes
$f(x)\geq\ln x,$ (37)
supplemented by $f(1)=0$. Our choice is $f(x)=x-1$. The choice made by Busiello, Gupta, and Maritan busiello2020 is
$\Psi_{21}=W_{21}P_{1}\ln\frac{P_{2}}{P_{1}}.$ (38)
and corresponds to $f(x)=\ln x$. But this expression for $\Psi_{21}$ makes the corresponding entropy production $\Pi_{21}=0$ to vanish identically.
An essential property required for the production of entropy is that it vanishes at thermodynamic equilibrium. As the production of entropy is composed of a sum of terms that are each one nonnegative then each term should vanish. In the case of the Schnakenberg formula (2), the vanishing of each term implies that the transition rates should fulfill the condition
$W_{ij}P_{j}^{0}=W_{ji}P_{i}^{0},$ (39)
where $P_{i}^{0}$ is the equilibrium distribution. In the case of formula (4), the condition is
$P_{j}^{0}=P_{i}^{0}.$ (40)
Therefore if both types of transitions are present, unidirectional and bidirectional, then the thermodynamic equilibrium is possible if $W_{ij}=W_{ji}$.
We show next that the production of entropy is a minimum at the equilibrium. To this end we write $P_{i}=q_{i}+\delta p_{i}$ and replace in
$\Pi_{ij}=W_{ij}[q_{j}\ln\frac{q_{j}}{q_{i}}-(q_{j}-q_{i})].$ (41)
Up to second order in $\delta p_{i}$, we find
$\delta\Pi_{ij}=W_{ij}[\delta p_{j}\ln\frac{q_{j}}{q_{i}}-\delta p_{i}(\frac{q_{j}}{q_{i}}-1)]+\frac{1}{2}W_{ij}q_{j}(\frac{\delta p_{j}}{q_{j}}-\frac{\delta p_{i}}{q_{i}})^{2}.$ (42)
If $q_{i}=P_{i}^{0}$ the terms linear in $\delta p_{i}$ vanish and we are left with
$\delta\Pi_{ij}=\frac{W_{ij}}{2P_{i}^{0}}(\delta p_{j}-\delta p_{i})^{2},$ (43)
which shows that the entropy production is a minimum at equilibirum. The value of $\Pi_{ij}$ at equilibrium also vanish as expected.
We remark that the additivity of the entropy flux follows immediately from the expression (22) and is implicit is our numerical calculation. Notice that the plots in figure 1 shows the entropy flux per site of the lattice.
VI Conclusion
We have proposed an expression for the rate of entropy production for the case in which the reverse transition rate vanishes but the forward transition rate is nonzero. This expression is positive definite and the corresponding entropy flux is linear in the probability distribution allowing its calculation as an average over this distribution. This expression allows us to determine the rate of entropy production for stochastic dynamic systems whose some of the transitions do not have the reverse. This is the case of models with with absorbing states such as the contact process.
We have applied our expression to the one-dimensional contact model and determined is rate of entropy production at the stationary state. We found that the rate of entropy per site is finite having a singularity at the critical point. The derivative of the production of entropy with respect to the annihilation parameter is also singular at the critical point, diverging at this point. The critical exponent associated to critical behavior of the rate of entropy production is found to be numerically equal to the exponent $\beta$ associated to the critical behavior of the particle density of the one-dimensional contact model.